# Asymptotic Analysis of High Order IMEX-RK Methods for ES-BGK Model at Navier-Stokes level

### v2

@@ 1 Introduction
where $f(t,x,v)$ is the distribution function which depends on time $t>0$, on the position of particles $x\in\mathbb{R}^{d_{x}}$ and on their velocity $v\in\mathbb{R}^{d_{v}}$. Here $\varepsilon$ is the so-called Knudsen number, defined as the ratio of the mean free path of molecules and the characteristic length scale of the physical problem. The dimension of space and velocity domains are denoted by $d_{x}$ and $d_{v}$, respectively.
@@
    $\mathcal{M}[f](t,x,v)=\frac{\rho(t,x)}{(2\pi T(t,x))^{d_{v}/2}}\exp\left(-\frac{|v-u(t,x)|^{2}}{2T(t,x)}\right)$ (1.4)
    where $\rho$, $u$, and $E$ are density, mean velocity and energy associated to $f$:
@@
where $\tau$ is the collision frequency that depends on $\rho$ and $T$. The BGK model replaces the Boltzmann collision operator with a simple relaxation toward the local Maxwellian. Note that the BGK model still maintains the conservation of mass, momentum, and energy, as well as the entropy inequality [13]. In addition, the BGK model describes the correct fluid limit as $\varepsilon\to 0$, i.e., at the leading order term $\varepsilon=0$, it yields the compressible Euler equations (see [13, 19]). Unfortunately, at the first-order correction in $\varepsilon$, the transport coefficients obtained at the Navier–Stokes level are not satisfactory. In particular, the Prandtl number defined by $\textrm{Pr}=\frac{\gamma}{\gamma-1}\frac{\mu}{\kappa},$ which relates the viscosity $\mu$ to the heat conductivity $\kappa$ of gases is fixed by $1$. Here, the polytropic constant $\gamma$ for monatomic molecule with transional motions is given by $\gamma=\frac{d_{v}+2}{d_{v}}$. Note that for most realistic gases, we have $Pr<1$. In addition, in the hard-sphere model for monoatomic gases ($\gamma=5/3$) in Boltzmann equation, its Prandtl number is very close to $2/3$.
Many variants of the BGK model have been proposed in order to give the correct transport coefficeints at the Navier–Stokes level. In [20], Holway proposed a model that possesses several desirable properties. It not only satisfies the correct conservation laws and the entropy condition, but also yields the Navier-Stokes equations with a Prandtl number less than one through the Chapman-Enskog expansion. This model is known as the ellipsoidal statistical model (ES-BGK) and reads:
@@
respectively, and $Id$ is the $d_{v}\times d_{v}$ identity matrix.
@@
$\partial_{t}\left(\begin{array}[]{l}\rho\\ \rho u\\ E\end{array}\right)+\nabla_{x}\left(\begin{array}[]{l}\rho u\\ \rho u\otimes u+pId\\ (E+p)u\end{array}\right)=\left(\begin{array}[]{l}0\\ \varepsilon\nabla_{x}\cdot(\mu\sigma(u))\\ \varepsilon\nabla_{x}\cdot(\mu\sigma(u)u+q)\end{array}\right)$ (1.21)
@@
and this implies that desired such number can be recovered by choosing $\nu$ appropriately. Note that the equation for the total energy can be replaced by the equation for the temperature $T$, [2]
@@
On the construction of AP and AA methods for kinetic models, IMEX time discretization [1, 23, 24, 11, 5, 6] have been successfully applied and studied.
We focus our literature reviews to the works on the NS limit based on IMEX time discretization applied to BGK or ES-BGK model. There is, of course, a considerable amount of literature on the use of IMEX-RK methods for BGK-type equations, (see, e.g.,[17, 16, 9]). For instance, in [3, 26], the authors considered a micro-macro decomposition of the BGK equation and then applied IMEX-RK schemes to the resulting coupled system. In [18], authors introduced a first order IMEX method for ES-BGK model and showed that it is consistent to first order time discretization of CNS equations. High order IMEX-RK methods of type CK (or type II) [21] and IMEX multistep methods [17] are also similarly applied to ES-BGK model, and the methods are shown to be able to capture the NS limit under suitable conditions. However, reference [21] focuses solely on a specific subclass of IMEX schemes (namely, type CK or type II methods, [9]).
@@
Another novelty of the present work is that we numerically demonstrate the uniform accuracy of IMEX-RK methods applied to ES-BGK model for a wide range of Knudsen numbers presented in [7]. These IMEX-RK schemes were originally developed and analyzed for a specific class of hyperbolic relaxation systems. In the context of the ES-BGK model, we derive additional order conditions (see Theorem 3.8) that ensure uniform accuracy for schemes of type II. We show that, under suitable assumptions on the coefficients of such schemes, the considered schemes satisfy these additional conditions and maintain the order of accuracy throughout the entire range of Knudsen numbers. These conditions are sufficient but not necessary. However, the price to pay is that the IMEX-RK schemes proposed in [7] require more stages than those commonly used.
@@ 2 IMEX-RK schemes for ES-BGK equation
The first condition in Definition 2.3 guarantees that an $A$-stable implicit tableau is $L$-stable [^1] A Runge-Kutta method is $L$-stable if $\lim_{z\to\infty}R(z)=0$, where R(z)R(z)is the stability function of the method. If the $A$-stable RK-scheme is SA, i.e., $b^{\top}A^{-1}{\bf e}=1$, it follows that $R(\infty)=\lim_{z\to\infty}R(z)=1-b^{\top}A^{-1}{\bf e}=0$, [25]. Here ${\bf e}=(1,1,...,1)^{\top}$ is a unit vector of length $s$.. The AP property of IMEX-RK schemes are strongly related to the $L$-stability of the implicit part of the scheme. Finally, if the IMEX-RK scheme is GSA we have $f^{n+1}=f^{(s)}$, i.e., the numerical solution coincides with the last stage of the method.
@@
    where ${\bf F}=(f^{(1)},f^{(2)},...,f^{(s)})^{\top}$, $L({\bf F})=(L(f^{(1)}),...,L(f^{(s)}))^{\top}$, being $L(f^{(k)})=v\cdot\nabla_{x}f^{(k)}$ for all k=1,,sk=1,...,s, $\mathcal{G}[{\bf F}]:=(\mathcal{G}(f^{(1)}),...,\mathcal{G}(f^{(s)}))^{\top}$, $\bar{{\bf\tau}}:=diag(\tau^{(1)},...,\tau^{(s)})$ is a diagonal matrix, and ${\bf e}=(1,1,...,1)^{\top}$ is a unit vector of length $s$.
@@
    with $\hat{{\bf e}}=(1,1,...,1)^{\top}$ a unit vector of length s1s-1. From the second equation in (2.30) we get
    $\Delta t\hat{\bar{\tau}}(\mathcal{G}[\hat{{\bf F}}]-\hat{{\bf F}})=\varepsilon\hat{A}^{-1}\left[\hat{\bf F}-f^{n}\hat{\bf e}+\Delta t\tilde{a}L(f^{n})+\Delta t\hat{\tilde{A}}L(\hat{\bf F})\right]-\Delta t\hat{A}^{-1}a{\tau^{n}}(\mathcal{G}[f^{n}]-f^{n}).$ (2.31)
@@ 3 Asymptotic Properties of the IMEX RK Schemes
Note that using (3.36) we can obtain $\rho^{(k)}$, $u^{(k)}$ and $T^{(k)}$ at every stage $k$, which enable us to compute explicitly $\tau^{(k)}$. To evaluate $\mathcal{G}[f^{(k)}]$, however, we need to determine the stress tensor $\Theta^{(k)}$ in (1.10). To achieve this, we first define the tensor $\Sigma$ as follows
@@
$\Sigma^{*}=\Sigma^{n}-\Delta t\sum_{\ell=1}^{k-1}\tilde{a}_{k,\ell}\nabla_{x}\biggl\langle v\otimes v\,vf^{(\ell)}\biggr\rangle+\frac{\Delta t}{\varepsilon}\sum_{\ell=1}^{k-1}a_{k\ell}(1-\nu)\tau^{(\ell)}[\rho^{(\ell)}(T^{(\ell)}Id+u^{(\ell)}\otimes u^{(\ell)})-\Sigma^{(\ell)}].$
@@
$\begin{split}\Sigma^{(k)}&=\frac{\varepsilon}{\varepsilon+(1-\nu)\tau^{(k)}\Delta ta_{kk}}\Sigma^{*}+\frac{\Delta t(1-\nu)\tau^{(k)}a_{kk}}{\varepsilon+(1-\nu)\tau^{(k)}\Delta ta_{kk}}\rho^{(k)}\left({T}^{(k)}Id+u^{(k)}\otimes u^{(k)})\right),\\ f^{(k)}&=\frac{\varepsilon}{\varepsilon+\tau^{(k)}\Delta ta_{kk}}f_{*}^{(k)}+\frac{\Delta t\tau^{(k)}a_{kk}}{\varepsilon+\tau^{(k)}\Delta ta_{kk}}\mathcal{G}[f^{(k)}],\end{split}$ (3.42)
@@
$\begin{split}\Sigma^{n+1}&=\Sigma^{n}-\Delta t\sum_{k=1}^{s}\tilde{b}_{k}\nabla_{x}\biggl\langle v\otimes v\,vf^{(k)}\biggr\rangle+\frac{\Delta t}{\varepsilon}\sum_{k=1}^{s}{b}_{k}\biggl\langle v\otimes v\,\left(\tau^{(k)}\left(\mathcal{G}[f^{(k)}]-f^{(k)}\right)\right)\biggr\rangle,\\ f^{n+1}&=f^{n}-\Delta t\sum_{k=1}^{s}\tilde{b}_{k}v\cdot\nabla_{x}f^{(k)}+\frac{\Delta t}{\varepsilon}\sum_{k=1}^{s}{b}_{k}\tau^{(k)}\left(\mathcal{G}[f^{(k)}]-f^{(k)}\right).\end{split}$ (3.43)
@@
Corollary 3.2 claims that an important property of the IMEX schemes of type I is obtained if in the limit $\varepsilon\to 0$ the distribution function is projected over the equilibrium, i.e. $f^{n+1}=\mathcal{M}(f^{n+1})$. From (2.29) it is clear that this property is achieved if the scheme is GSA i.e., $b^{\top}A^{-1}={\bf e}^{\top}_{s}$, with ${\bf e}^{\top}_{s}=(0,...0,1)$ vector of length $s$, and $b^{\top}A^{-1}\tilde{A}=\tilde{b}^{\top}$, [9, 16].
@@
$\begin{split}\rho^{(k)}&=\rho^{n}-\Delta t\sum_{\ell=1}^{k-1}\tilde{a}_{k\ell}\textrm{div}_{x}(\rho^{(\ell)}u^{(\ell)}),\\ (\rho u)^{(k)}&=(\rho u)^{n}-\Delta t\sum_{\ell=1}^{k-1}\tilde{a}_{k\ell}\textrm{div}_{x}((\rho u)^{(\ell)}\otimes u^{(\ell)}+\rho^{(\ell)}T^{(\ell)}I),\\ E^{(k)}&=E^{n}-\Delta t\sum_{\ell=1}^{k-1}\tilde{a}_{k\ell}\textrm{div}_{x}([E^{(\ell)}+\rho^{(\ell)}T^{(\ell)}]u^{(\ell)}),\end{split}$ (3.45)
@@
$\begin{split}\rho^{n+1}&=\rho^{n}-\Delta t\sum_{k=1}^{s}\tilde{b}_{k}\nabla_{x}(\rho^{(k)}u^{(k)}),\\ (\rho u)^{n+1}&=(\rho u)^{n}-\Delta t\sum_{k=1}^{s}\tilde{b}_{k}\nabla_{x}((\rho u)^{(k)}\otimes u^{(k)}+\rho^{(k)}T^{(k)}I),\\ E^{n+1}&=E^{n}-\Delta t\sum_{k=1}^{s}\tilde{b}_{k}\nabla_{x}([E^{(k)}+\rho^{(k)}T^{(k)}]u^{(k)}).\end{split}$ (3.46)
@@
$T^{(k)}=T^{n}-\Delta t\sum_{\ell=1}^{k-1}\tilde{a}_{k\ell}\left(u^{(\ell)}\cdot\nabla_{x}T^{(\ell)}+\frac{2}{d_{v}}T^{(\ell)}\textrm{div}_{x}u^{(\ell)}\right),$ (3.47)
$T^{n+1}=T^{n}-\Delta t\sum_{k=1}^{s}\tilde{b}_{k}\left(u^{(k)}\cdot\nabla_{x}T^{(k)}+\frac{2}{d_{v}}T^{(k)}\textrm{div}_{x}u^{(k)}\right).$ (3.48)
@@
In this section, we analyze the asymptotic behavior of IMEX-RK schemes of type I and II for ES-BGK equations (1.7), and prove that for small value of $\varepsilon$, these schemes asymptotically capture the NS limit without resolving $\varepsilon$, providing a numerical scheme for the corresponding CNS equations (1.21). At the Navier-Stokes level, (for small but non-negligible Knudsen numbers, $\varepsilon\ll 1$), the GSA condition is not crucial to ensure consistency with CNS equations for both types of IMEX-RK schemes. Therefore, for the subsequent analysis, we do not impose the GSA condition on the scheme.
IMEX-RK scheme of type I. We first analyze the IMEX-RK method of type I.
@@
$\langle\phi{\mathcal{M}}[{\bf F}]\rangle=\langle\phi\mathcal{M}[f^{n}]{\bf e}\rangle-\Delta t\tilde{A}\langle\phi L(\mathcal{M}[{\bf F}])\rangle-\varepsilon\Delta t\tilde{A}\nabla_{x}\cdot\langle v\phi{\bf f}_{1}\rangle,$
and, defining $\nabla_{x}F({\bf U})=\langle\phi L(\mathcal{M}[{\bf F}])\rangle$ the flux vector of ${\bf U}=\langle\phi\mathcal{M}[{\bf F}]\rangle$, it follows
${\bf U}=U^{n}{\bf e}-\Delta t\tilde{A}\nabla_{x}F({\bf U})-\varepsilon\Delta t\tilde{A}\nabla_{x}\langle v\phi{\bf f}_{1}\rangle.$ (3.50)
@@
${\bf U}=U^{n}{\bf e}-\Delta t\tilde{A}\nabla_{x}F({\bf U})-\varepsilon\Delta t\tilde{A}\nabla_{x}H({\bf U}),$ (3.51)

### v3

@@ 1 Introduction
where $f(t,x,v)$ is the distribution function which depends on time $t>0$, on the position of particles $x=(x_{1},\cdots,x_{d_{x}})\in\mathbb{R}^{d_{x}}$ and on their velocity $v=(v_{1},\cdots,v_{d_{v}})\in\mathbb{R}^{d_{v}}$. Here $\varepsilon$ is the so-called Knudsen number, defined as the ratio of the mean free path of molecules and the characteristic length scale of the physical problem. The dimension of space and velocity domains are denoted by $d_{x}$ and $d_{v}$, respectively.
@@
    $\mathcal{M}[f](t,x,v)=\frac{\rho(t,x)}{(2\pi T(t,x))^{d_{v}/2}}\exp\left(-\frac{|v-u(t,x)|^{2}}{2T(t,x)}\right),$ (1.4)
    where $\rho$, $u=(u_{1},\cdots,u_{d_{v}})$, and $E$ are density, mean velocity and energy associated to $f$:
@@
where $\tau$ is the collision frequency that depends on $\rho$ and $T$. The BGK model replaces the Boltzmann collision operator with a simple relaxation toward the local Maxwellian. Note that the BGK model still maintains the conservation of mass, momentum, and energy, as well as the entropy inequality [13]. In addition, the BGK model describes the correct fluid limit as $\varepsilon\to 0$, i.e., at the leading order term $\varepsilon=0$, it yields the compressible Euler equations (see [13, 19]). Unfortunately, at the first-order correction in $\varepsilon$, the transport coefficients obtained at the Navier–Stokes level are not satisfactory. In particular, the Prandtl number defined by $\textrm{Pr}=\frac{\gamma}{\gamma-1}\frac{\mu}{\kappa},$ which relates the viscosity $\mu$ to the heat conductivity $\kappa$ of gases is fixed by $1$. Here, the polytropic constant $\gamma$ for monatomic molecule with transional motions is given by $\gamma=\frac{d_{v}+2}{d_{v}}$. Note that for most realistic gases, we have Pr<1<1. In addition, in the hard-sphere model for monoatomic gases ($\gamma=5/3$) in Boltzmann equation, its Prandtl number is very close to $2/3$.
Many variants of the BGK model have been proposed in order to give the correct transport coefficients at the Navier–Stokes level. In [20], Holway proposed a model that possesses several desirable properties. It not only satisfies the correct conservation laws and the entropy condition, but also yields the Navier-Stokes equations with a Prandtl number less than one through the Chapman-Enskog expansion. This model is known as the ellipsoidal statistical model (ES-BGK) and reads:
@@
respectively, where $Id$ is the $d_{v}\times d_{v}$ identity matrix and $(v-u)\otimes(v-u)$ denotes the tensor product [^1] The tensor product of two vectors $a=(a_{1},...,a_{d}),b=(b_{1},...,b_{d})\in\mathbb{R}^{d}$ means the outer product of $u$ and $v$, i.e., $a\otimes b\in\mathbb{R}^{d\times d}$ with $(a\otimes b)_{ij}=a_{i}b_{j}$. of two vectors (vu)(v-u) and (vu)(v-u).
@@
$\partial_{t}\left(\begin{array}[]{l}\rho\\ \rho u\\ E\end{array}\right)+\nabla_{x}\cdot\left(\begin{array}[]{l}\rho u\\ \rho u\otimes u+pId\\ (E+p)u\end{array}\right)=\left(\begin{array}[]{l}0\\ \varepsilon\nabla_{x}\cdot(\mu\sigma(u))\\ \varepsilon\nabla_{x}\cdot(\mu\sigma(u)u+q)\end{array}\right)$ (1.21)
@@
and this implies that desired such number can be recovered by choosing $\nu$ appropriately. Note that the equation for the total energy can be replaced by the equation for the temperature $T$ [2] as follows:
@@
On the construction of AP and AA methods for kinetic models, IMEX time discretization [1, 23, 24, 9, 11, 12] have been successfully applied and studied. We focus our literature reviews to the works on the NS limit based on IMEX time discretization applied to BGK or ES-BGK model. There is, of course, a considerable amount of literature on the use of IMEX-RK methods for BGK-type equations, (see, e.g.,[17, 16, 7]). For instance, in [3, 26], the authors considered a micro-macro decomposition of the BGK equation and then applied IMEX-RK schemes to the resulting coupled system. In [18], authors introduced a first order IMEX method for ES-BGK model and showed that it is consistent to first order time discretization of CNS equations. High order IMEX-RK methods of type CK (or type II) [21] and IMEX multistep methods [17] are also similarly applied to ES-BGK model, and the methods are shown to be able to capture the NS limit under suitable conditions. However, reference [21] focuses solely on a specific subclass of IMEX schemes (namely, type CK or type II methods, [7]).
@@
Another novelty of the present work is that we show numerically the uniform accuracy of specific type of IMEX-RK methods, originally introduced in [8] when applied to ES-BGK model for a wide range of Knudsen numbers. These IMEX-RK schemes were originally developed and analyzed for a specific class of hyperbolic relaxation systems. In the context of the ES-BGK model, we show that, under suitable assumptions on the coefficients of such schemes introduced in [8], consistency with CNS equations is ensured (see Theorem 3.8). Moreover, additionally order conditions allow the schemes to maintain the order of accuracy throughout the entire range of Knudsen numbers, thereby guaranteeing uniform accuracy. These conditions are sufficient but not necessary. However, the price to pay is that the IMEX-RK schemes proposed in [8] require more stages than those commonly used.
@@ 2 IMEX-RK schemes for ES-BGK equation
The first condition in Definition 2.3 guarantees that an $A$-stable implicit tableau is $L$-stable [^2] A Runge-Kutta method is $L$-stable if $\lim_{z\to\infty}R(z)=0$, where $R(z)$ is the stability function of the method. If the $A$-stable RK-scheme is SA, i.e., $b^{\top}A^{-1}{\bf e}=1$, it follows that $R(\infty)=\lim_{z\to\infty}R(z)=1-b^{\top}A^{-1}{\bf e}=0$, [25]. Here ${\bf e}=(1,1,...,1)^{\top}$ is a vector of length $s$.. The AP property of IMEX-RK schemes are strongly related to the $L$-stability of the implicit part of the scheme. Finally, if the IMEX-RK scheme is GSA we have $f^{n+1}=f^{(s)}$, i.e., the numerical solution coincides with the last stage of the method.
@@
    where ${\bf F}=(f^{(1)},f^{(2)},...,f^{(s)})^{\top}\in\mathbb{R}^{s}$, $L({\bf F})=(L(f^{(1)}),...,L(f^{(s)}))^{\top}\in\mathbb{R}^{s}$, being $L(f^{(k)})=v\cdot\nabla_{x}f^{(k)}$ for all k=1,,sk=1,...,s, $\mathcal{G}[{\bf F}]:=(\mathcal{G}(f^{(1)}),...,\mathcal{G}(f^{(s)}))^{\top}\in\mathbb{R}^{s}$. A diagonal matrix $\bar{{\bf\tau}}:=diag(\tau^{(1)},...,\tau^{(s)})$ is defined with the relaxation time $\tau^{(i)}$ associated to $f^{(i)}$ and ${\bf e}=(1,1,...,1)^{\top}$ is a vector of length $s$.
@@
    where $\hat{\bf F}=(f^{(2)},...,f^{(s)})^{\top}\in\mathbb{R}^{s-1}$, $L(\hat{\bf F})=(L(f^{(2)}),...,L(f^{(s)}))^{\top}\in\mathbb{R}^{s-1}$, $\hat{\bf e}=(1,1,...,1)^{\top}\in\mathbb{R}^{s-1}$, $\hat{\tilde{b}}=(\tilde{b}_{2},\tilde{b}_{3},...,\tilde{b}_{s})^{\top}$, $\tau^{n}=\tau^{(1)}$, $\hat{\bar{\tau}}:=diag(\tau^{(2)},...,\tau^{(s)})$ and $\mathcal{G}[{\bf\hat{F}}]:=(\mathcal{G}(f^{(2)}),...,\mathcal{G}(f^{(s)}))^{\top}\in\mathbb{R}^{s-1}$. From the second equation in (2.30) we get
    $\Delta t\hat{\bar{\tau}}(\mathcal{G}[\hat{{\bf F}}]-\hat{{\bf F}})=\varepsilon\hat{A}^{-1}\left[\hat{\bf F}-f^{n}\hat{\bf e}+\Delta t\tilde{a}L(f^{n})+\Delta t\hat{\tilde{A}}L(\hat{\bf F})\right]-\Delta t\hat{A}^{-1}a{\tau^{n}}(\mathcal{G}[f^{n}]-f^{n}),$ (2.31)
@@ 3 Asymptotic Properties of the IMEX RK Schemes
Note that using (3.36) we can obtain $\rho^{(k)}$, $u^{(k)}$ and $T^{(k)}$ at every stage $k$, which enables us to compute explicitly $\tau^{(k)}$. To evaluate $\mathcal{G}[f^{(k)}]$, however, we need to determine the stress tensor $\Theta^{(k)}$ in (1.10). To achieve this, we first define the tensor $\Sigma$ as follows
@@
$\Sigma^{*}=\Sigma^{n}-\Delta t\sum_{\ell=1}^{k-1}\tilde{a}_{k,\ell}\nabla_{x}\cdot\biggl\langle v\otimes v\,vf^{(\ell)}\biggr\rangle+\frac{\Delta t}{\varepsilon}\sum_{\ell=1}^{k-1}a_{k\ell}(1-\nu)\tau^{(\ell)}[\rho^{(\ell)}(T^{(\ell)}Id+u^{(\ell)}\otimes u^{(\ell)})-\Sigma^{(\ell)}].$
@@
$\begin{split}\Sigma^{(k)}&=\frac{\varepsilon}{\varepsilon+(1-\nu)\tau^{(k)}\Delta ta_{kk}}\Sigma^{*}+\frac{\Delta t(1-\nu)\tau^{(k)}a_{kk}}{\varepsilon+(1-\nu)\tau^{(k)}\Delta ta_{kk}}\rho^{(k)}\left({T}^{(k)}Id+u^{(k)}\otimes u^{(k)}\right),\\ f^{(k)}&=\frac{\varepsilon}{\varepsilon+\tau^{(k)}\Delta ta_{kk}}f_{*}^{(k)}+\frac{\Delta t\tau^{(k)}a_{kk}}{\varepsilon+\tau^{(k)}\Delta ta_{kk}}\mathcal{G}[f^{(k)}],\end{split}$ (3.42)
@@
$\begin{split}\Sigma^{n+1}&=\Sigma^{n}-\Delta t\sum_{k=1}^{s}\tilde{b}_{k}\nabla_{x}\cdot\biggl\langle v\otimes v\,vf^{(k)}\biggr\rangle+\frac{\Delta t}{\varepsilon}\sum_{k=1}^{s}{b}_{k}\biggl\langle v\otimes v\,\left(\tau^{(k)}\left(\mathcal{G}[f^{(k)}]-f^{(k)}\right)\right)\biggr\rangle,\\ f^{n+1}&=f^{n}-\Delta t\sum_{k=1}^{s}\tilde{b}_{k}v\cdot\nabla_{x}f^{(k)}+\frac{\Delta t}{\varepsilon}\sum_{k=1}^{s}{b}_{k}\tau^{(k)}\left(\mathcal{G}[f^{(k)}]-f^{(k)}\right).\end{split}$ (3.43)
@@
Corollary 3.2 claims that an important property of the IMEX schemes of type I is obtained if in the limit $\varepsilon\to 0$ the distribution function is projected over the equilibrium, i.e., $f^{n+1}=\mathcal{M}(f^{n+1})$. From (2.29) it is clear that this property is achieved if the scheme is GSA i.e., $b^{\top}A^{-1}={\bf e}^{\top}_{s}$, with ${\bf e}^{\top}_{s}=(0,...0,1)$ vector of length $s$, and $b^{\top}A^{-1}\tilde{A}=\tilde{b}^{\top}$, [7, 16].
@@
$\begin{split}\rho^{(k)}&=\rho^{n}-\Delta t\sum_{\ell=1}^{k-1}\tilde{a}_{k\ell}\nabla_{x}\cdot(\rho^{(\ell)}u^{(\ell)}),\\ (\rho u)^{(k)}&=(\rho u)^{n}-\Delta t\sum_{\ell=1}^{k-1}\tilde{a}_{k\ell}\nabla_{x}\cdot\left((\rho u)^{(\ell)}\otimes u^{(\ell)}+\rho^{(\ell)}T^{(\ell)}Id\right),\\ E^{(k)}&=E^{n}-\Delta t\sum_{\ell=1}^{k-1}\tilde{a}_{k\ell}\nabla_{x}\cdot\left((E^{(\ell)}+\rho^{(\ell)}T^{(\ell)})u^{(\ell)}\right),\end{split}$ (3.45)
@@
$\begin{split}\rho^{n+1}&=\rho^{n}-\Delta t\sum_{k=1}^{s}\tilde{b}_{k}\nabla_{x}\cdot(\rho^{(k)}u^{(k)}),\\ (\rho u)^{n+1}&=(\rho u)^{n}-\Delta t\sum_{k=1}^{s}\tilde{b}_{k}\nabla_{x}\cdot\left((\rho u)^{(k)}\otimes u^{(k)}+\rho^{(k)}T^{(k)}Id\right),\\ E^{n+1}&=E^{n}-\Delta t\sum_{k=1}^{s}\tilde{b}_{k}\nabla_{x}\cdot\left((E^{(k)}+\rho^{(k)}T^{(k)})u^{(k)}\right).\end{split}$ (3.46)
@@
$T^{(k)}=T^{n}-\Delta t\sum_{\ell=1}^{k-1}\tilde{a}_{k\ell}\left(u^{(\ell)}\cdot\nabla_{x}T^{(\ell)}+\frac{2}{d_{v}}T^{(\ell)}\nabla_{x}\cdot u^{(\ell)}\right),$ (3.47)
$T^{n+1}=T^{n}-\Delta t\sum_{k=1}^{s}\tilde{b}_{k}\left(u^{(k)}\cdot\nabla_{x}T^{(k)}+\frac{2}{d_{v}}T^{(k)}\nabla_{x}\cdot u^{(k)}\right).$ (3.48)
@@
In this section, we analyze the asymptotic behavior of IMEX-RK schemes of type I and II for ES-BGK equations (1.7), and prove that for small value of $\varepsilon$, these schemes asymptotically capture the NS limit without resolving $\varepsilon$, providing a numerical scheme for the corresponding CNS equations (1.21). The two main Theorems below not only extend the asymptotic analysis of [21] by considering genereal IMEX RK methods, but also result in explicit-type RK schemes for CNS equations which are consistent with the results of [8].
At the Navier-Stokes level, (for small but non-negligible Knudsen numbers, $\varepsilon\ll 1$), the GSA condition is not crucial to ensure consistency with CNS equations for both types of IMEX-RK schemes. Therefore, for the subsequent analysis, we do not impose the GSA condition on the scheme.
$\bullet$ IMEX-RK scheme of type I. We first analyze the IMEX-RK method of type I. The following shows that the result in [8] can be generalized to ES-BGK model.
@@
$B=\tilde{A}A^{-1}\tilde{A},\quad\omega^{\top}=\tilde{b}^{T}A^{-1}\tilde{A}$.
@@
$\langle\phi{\mathcal{M}}[{\bf F}]\rangle=\langle\phi\mathcal{M}[f^{n}]{\bf e}\rangle-\Delta t{\bf\tilde{A}}\langle\phi L(\mathcal{M}[{\bf F}])\rangle-\varepsilon\Delta t{\bf\tilde{A}}\nabla_{x}\cdot\langle v\phi{\bf f}_{1}\rangle,$
where ${\bf\tilde{A}}:=\tilde{A}\otimes_{K}I_{d_{v}+2}\in\mathbb{R}^{(2+d_{v})s\times(2+d_{v})s}$, $I_{d_{v}+2}$ is the $(d_{v}+2)\times(d_{v}+2)$ identity matrix, and the symbol $\otimes_{K}$ denotes the Kronecker product [^4] We use the symbol $\otimes_{K}$ to denote the Kronecker product: If $A$ is a $m\times n$ matrix and $B$ is a $p\times q$ matrix, then the Kronecker product of $A,B$ is a $mp\times nq$ matrix that can be written in block form as $A\otimes_{K}B=\begin{pmatrix}a_{11}B&\cdots&a_{1n}B\\ \vdots&&\vdots\\ a_{m1}B&\cdots&a_{mn}B\\ \end{pmatrix}$. Defining the flux vector $\nabla_{x}\cdot F(U^{(i)})=\langle\phi L(\mathcal{M}[F^{(i)}])\rangle$ with $U^{(i)}=\langle\phi\mathcal{M}[F^{(i)}]\rangle$, it follows
${\bf U}={\bf e}\otimes_{K}U^{n}-\Delta t{\bf\tilde{A}}\nabla_{x}\cdot F({\bf U})-\varepsilon\Delta t{\bf\tilde{A}}\nabla_{x}\cdot\langle v\phi{\bf f}_{1}\rangle,$ (3.50)
where ${\bf e}=(1,1,...,1)^{\top}\in\mathbb{R}^{s}$,
$\begin{split}{\bf U}&=\begin{pmatrix}U^{(1)},U^{(2)},\cdots,U^{(s)}\end{pmatrix}^{\top}\in\mathbb{R}^{(2+d_{v})s},\\ U^{(i)}&=(\rho^{(i)},\rho^{(i)}u^{(i)},E^{(i)})^{\top}\in\mathbb{R}^{2+d_{v}}\\ \nabla_{x}\cdot F({\bf U})&=\left(\nabla_{x}\cdot F(U^{(1)}),\nabla_{x}\cdot F(U^{(2)}),...,\nabla_{x}\cdot F(U^{(s)})\right)^{\top}\in\mathbb{R}^{(2+d_{v})s},\\ \nabla_{x}\cdot F(U^{(i)})&=\left(\begin{array}[]{c}\nabla_{x}\cdot(\rho^{(i)}u^{(i)})\\ \nabla_{x}\cdot(\rho^{(i)}u^{(i)}\otimes u^{(i)}+\rho^{(i)}T^{(i)}Id)\\ \nabla_{x}\cdot((E^{(i)}+\rho^{(i)}T^{(i)})u^{(i)})\end{array}\right)\in\mathbb{R}^{2+d_{v}}\,(i=1,...,s).\end{split}$ (3.51)
@@
${\bf U}={{\bf e}\otimes_{K}U^{n}}-\Delta t{\bf\tilde{A}}\nabla_{x}\cdot F({\bf U})-\varepsilon\Delta t{\bf\tilde{A}}\nabla_{x}\cdot{\bf H}({\bf U}),$ (3.52)